Could someone please help me?
Given f(x)=x^2 and g(x)=3x-1, find (f*g)(2) and (g*f)(x)
Any help will be greatly appreciated.

Answer by stanbon(75887)

Given f(x)=x^2 and g(x)=3x-1, find (fog)(2) and (gof)(x)
-------------
(fog)(2) = f[g(2)] = f[3*2-1] = f[5] = 25
---------------
(gof)(x) = g[f(x)] = g[x^2] = 3x^2-1
==================
Cheers,
Stan H.